Disk VSN: HOB+0120/32000/1024 | Data volume at beginning: 2590.634 GB
- 1800 UT Experiment started ok (Jay).
- 1014 UT Issue with dbbc, schedule stopped. Last scan 2018-453 (Eloise)
Disk VSN: HOB+0131/32000/1024 | Data volume at beginning: 7080.238 GB
- 1800UT Experiment started ok (Jay).
- 1115UT - 1155UT Antenna windstowed, then stuck. Missed scans 1448-648 through to 2348-432 (Eloise)